10 mL 0.15 M NaOH and 10 mL 0.1 M HCl are mixed to make a buffer. ka=1.8*10^-5 . what is the pH of this buffer?
1
Expert's answer
2015-04-15T13:43:15-0400
This solution is not a buffer because buffer is a solution which contains a weak acid and its conjugate base or a weak base of its conjugate acid.
The number of moles of NaOH is 0.01*0.15 = 0.0015 moles
The number o moles of HCI is 0.01*0.1 = 0.001 moles
The reactants react in the proportion 1:1, so the number of moles of NaOH which is left is 0.0015-0.001 = 0.0005 moles
The total volume of the solution is 20mL, or 0.02L.
So, the concentration of NaOH is 0.0005moles/0.02L = 0.025M
pOH = -log0.025 = 1.6
pH = 14 - 1.6 = 12.4